# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

### v1

@@ 6 Conclusion
In this work, we extended the Logarithmic Smoothing learning principle to the adaptive setting. We derived two novel adaptive algorithms for off-policy learning and established convergence guarantees to their respective optima. One of the algorithms enjoys accelerated convergence under mild assumptions and demonstrated improved empirical performance. These theoretical and practical benefits highlight the potential of combining online PAC-Bayes techniques with adaptive contextual bandits. Importantly, online PAC-Bayes methods do not rely on the i.i.d. assumption, opening promising directions for future research. In particular, relaxing the assumptions of a fixed context distribution and i.i.d. contexts would allow us to address distribution shift—where the context or cost distribution evolves during training. This setting better reflects real-world applications, where user behavior and responses naturally change over time.
Acknowledgments and Disclosure of Funding
M. Haddouche is supported by the European Union (ERC grant DYNASTY 101039676). The French government partly funded this work under the management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI).
Appendix A Limitations
This work develops theoretically grounded and practical learning approaches for the adaptive contextual bandit setting, where the decision-maker can dynamically adjust behavior policies to collect higher-quality data. Our theoretical analysis generalizes the static (batch) setting and highlights the advantages of multiple updates and deployments. While this represents a clear improvement over the batch setting, the proposed algorithm is still reliant on the initial behavior policy and may not converge to the optimal policy if the starting policy fails to explore optimal actions. Another limitation arises from the offline contextual bandit framework itself, which assumes i.i.d. observations. Although this assumption may be valid in simple scenarios, it becomes inadequate in real-world applications where user interests evolve over time. Extending our results to settings with distribution shifts, for example, could be achieved using similar proof techniques and would present an important avenue for future research.
Appendix B Broader impact statement
We contribute to the development of principled and practical pessimistic algorithms for the adaptive off-policy contextual bandit setting. These algorithms improve the robustness of decision-making by reducing unnecessary exploration, accelerating convergence to optimal policies, and prioritizing safety while mitigating uncertainty-related risks. By incorporating pessimistic strategies, our approach introduces a conservative bias that can lead to better outcomes in high-stakes environments where the cost of mistakes is significant. While our framework and algorithms have broad potential applicability, their social impact will ultimately depend on the specific domains in which they are deployed.
Appendix C Extended Related Work
@@
Off-Policy Learning in the contextual bandit setting, or counterfactual risk minimization [47] is a widely adopted framework for learning from past interactions. In most real world applications, large datasets summarizing past interactions are often available, allowing agents to improve their policies offline. In this setting, the goal is to find the optimal policy using previously collected interactions. Two primary approaches exist for learning from this data. The direct method [43] focuses on modeling the reward and subsequently derives a straightforward policy. This approach is well-understood, and the quality of the learned policy depends entirely on our ability to accurately model the reward. The other approach directly learns a policy from the data by minimizing an importance weighting estimator [15, 6] of the expected cost. The research community has predominantly focused on this second approach, as it presents significant theoretical challenges and offers greater generality by not relying on any reward modeling. At the core of directly learning a policy is the development of tailored off-policy estimators with improved concentration properties [6, 3, 35, 42], combined with tight generalization bounds allowing efficient learning [47, 41, 42]. The used methods are motivated from the pessimism principle [26] that was proven to be optimal for learning in these scenarios. Particularly, analyzing this learning problem from the PAC-Bayesian learning theory lens [31] proved efficient, improving the seminal work of [47]. This direction was further explored by [41, 3, 18, 42], reducing assumptions and improving the pessimistic principle, with latest developments [42] resulting in state of the art pessimistic approaches. We further develop this research direction and generalize it to the adaptive setting, where an online algorithm adapts the behavior policy to the environment to collect higher quality interactions.
Adaptive Off-Policy Learning.
Adaptive policy learning builds upon off-policy learning by enabling the decision-maker to refine the initial behavior policy and gather higher quality interactions with each new deployment. This approach is prevalent in practice as decision systems are continuously updated; however, its theoretical understanding remains limited and underexplored. Despite its practical significance, few works have addressed this question. [49] and subsequently [5] investigated learning from adaptively collected data in a purely online setting, where policies are updated after each step. In contrast to these online methods, our work considers updates occurring after the arrival of batches of arbitrary sizes. This better models real-world scenarios and provides a framework that interpolates between the fully static and fully online regimes. Our proposed algorithms achieve the convergence rate of the batch setting and accelerate it once we can refine our behavior policy. Furthermore, the analysis in [49, 5] relies on strong assumptions. Specifically, [49] employs a doubly robust estimator and provides regret guarantees but assumes uniform coverage to control variance. Under the same uniform coverage assumption, [5] introduces a novel maximal inequality and subsequently derives fast rate regret guarantees under a margin condition similar to ours. However, their results are limited to the online setting and also require uniform coverage, which can decay slowly over time. Recently, [48] extended Counterfactual Risk Minimization to the adaptive setting, allowing for intermediate policy deployments to facilitate the learning of better policies. Their algorithm achieves accelerated convergence rates under a Hölderian assumption but necessitates bounded importance weights, which, in the discrete action space case, is equivalent to the uniform coverage assumption. Additionally, their proposed algorithm only functions with exponentially increasing batch sizes, which is unrealistic, and only utilizes the last batch of interactions instead of the entire dataset. These limitations primarily stem from the theoretical tools employed, rendering the algorithm impractical. We analyze this problem from a PAC-Bayesian perspective and present the first algorithms that interpolate between full batch and full online regimes, matching the rates of both and operating under the coverage of the optimal actions alone.
@@
PAC-Bayes learning is a recent branch of learning theory which emerged in the late 90s via the seminal work of [45, 32, 33, 34] and later pursued by [7, 8]. Modern surveys are available to describe the various advances in the field [19, 1]. PAC-Bayes provide generalization bounds involving a complexity term (often a KL divergence), inspired here from the Bayesian learning paradigm of designing a posterior knowledge of the learning problem based on both training data and a prior knowledge of the considered situation. On the contrary, PAC-Bayes, while inspired from the Bayesian philosophy, relies historically on tools from information theory. This general approach benefits from additional flexibility as PAC-Bayes can be linked and applied to Bayesian learning (see 19) but also blurs the notion of prior and posterior distributions, now independent of the fundamental Bayes formula. This flexibility allowed PAC-Bayes to reach various sub-fields of learning theory: optimization dynamics of learning algorithms [36, 22, 2], reinforcement learning [17], online learning [30, 20], contrastive learning [37], generative models [9] or federated learning [40, 27].
Appendix D Proofs of Section 3
In this whole section, we assume that all datasets $S_{k}$ have the same cardinal $n_{k}=m$. This section contains the proof of Sec. 3.2 and Algorithm 1.
D.1 Proof of Theorem 3.2
@@ Appendix D Proofs of Section 3
Plugging this back into Eq. 4 yields, with probability at least $1-\delta$ over $S$, for any $k\geq 0$, any posterior $Q$:
$\sum_{j=0}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{k}}-\log\left(1-\lambda R(d_{\theta})\right)+\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}^{0}(a_{i}\mid x_{i})}\right)\right]\leq\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right).$
@@
$\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)=d_{\theta}(a,x)\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right).$
@@
D.2 Proof of Proposition 1
We prove the convergence guarantee of Algorithm 1.
@@ D.2 Proof of Proposition 1
We fix $k\geq 0$, an adaptive algorithm Alg and $P$ a data-free prior. To highlight the dependency on $\lambda$, we denote in this proof $\mathcal{S}_{j}=\mathcal{S}_{\lambda,j}$ Our goal is to apply the change of measure inequality on $\mathcal{H}$ to a specific function $f_{k}$. We define this function below, for any sample $S$ and any predictor $\theta$
@@
Plugging this back into Eq. 5 yields, with probability at least $1-\delta$ over $S$, for any $k\geq 0$, any $Q$:
@@
We also consider $\pi_{k+1}$ the output of Algorithm 1 time $k$ satisfying:
@@
Then using Sec. 3.2, and the definition of $\hat{Q}$ yields for any $\lambda$, data-free $P$ with probability at least $21-\delta/$ 2, for any $k$:
@@
Then, exploiting Eq. 6 with probability $21-\delta/$ 2 and taking a union bound yields, with probability at least $1-\delta$, for any $k$:

### v2

@@ 6 CONCLUSION
In this work, we extended the Logarithmic Smoothing learning principle to the sequential setting. We proposed two novel sequential algorithms for off-policy learning and established convergence guarantees to their respective optima. One of these algorithms enjoys accelerated convergence under mild assumptions and demonstrated improved empirical performance. These theoretical and practical benefits highlight the potential of combining online PAC-Bayes techniques with sequential contextual bandits. Our theoretical analysis requires our policy class to cover the optimal actions, thus weakening standard uniform coverage assumptions. An interesting open question is whether the structure of the pessimistic update can be exploited to relax this requirement. Moreover, in Appendix˜C, we showed that online PAC-Bayes techniques yield learning algorithms that relax the i.i.d. assumption on contexts, thereby addressing real-world scenarios with distribution shifts. A natural and promising direction is to investigate whether the (accelerated) convergence guarantees established here extend to this more general, non-stationary setting.
Acknowledgements
M.H. is partially supported by the European Research Council Starting Grant DYNASTY – 101039676. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] We provide all proofs and algorithm statement either in the main text or appendices.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] We study convergence guarantees of our algorithms as well as providing PAC-Bayes bound which justifies their existence.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] All results are stated with respect to our detailed theoretical setup of Sec.˜2.
    2. (b)
      Complete proofs of all theoretical results. [Yes] All proofs are detailed in the appendix due to space constraints.
    3. (c)
      Clear explanations of any assumptions. [Yes] We explain and contextualize our various assumptions in Secs.˜2, 3 and 4.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] Code lies in the supplementary material.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] All practical details are gathered in Secs.˜5 and E.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] Such details are retrievable in Secs.˜5 and E.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] All experiments were run on a single machine. Such details are retrievable in Appendix˜E.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] We mention all the inspirations we exploited as well as references of the datasets we used in Sec.˜5.
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable] We only exploited standard datasets such as MNIST and CIFAR100, publicly available and not requiring any specific consent to be used.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
APPENDIX
Appendix A EXTENDED RELATED WORK
@@
Off-Policy Learning in the contextual bandit setting, or counterfactual risk minimization [Swaminathan and Joachims, 2015] is a widely adopted framework for learning from past interactions. In most real world applications, large datasets summarizing past interactions are often available, allowing agents to improve their policies offline. In this setting, the goal is to find the optimal policy using previously collected interactions. Two primary approaches exist for learning from this data. The direct method [Sakhi et al., 2020] focuses on modeling the reward and subsequently derives a straightforward policy. This approach is well-understood, and the quality of the learned policy depends entirely on our ability to accurately model the reward. The other approach directly learns a policy from the data by minimizing an importance weighting estimator [Dudik et al., 2014, Bottou et al., 2013] of the expected cost. The research community has predominantly focused on this second approach, as it presents significant theoretical challenges and offers greater generality by not relying on any reward modeling. At the core of directly learning a policy is the development of tailored off-policy estimators with improved concentration properties [Bottou et al., 2013, Aouali et al., 2023, Metelli et al., 2021, Sakhi et al., 2024], combined with tight generalization bounds allowing efficient learning [Swaminathan and Joachims, 2015, Sakhi et al., 2023a, 2024]. The used methods are motivated by the pessimism principle [Jin et al., 2021] that was proven to be optimal for learning in these scenarios. Particularly, analyzing this learning problem from the PAC-Bayesian learning theory lens London and Sandler [2019] proved efficient, improving the seminal work of [Swaminathan and Joachims, 2015]. This direction was further explored by Sakhi et al. [2023a], Aouali et al. [2023], Gabbianelli et al. [2024], Sakhi et al. [2024], reducing assumptions and improving the pessimistic principle, with latest developments [Sakhi et al., 2024] resulting in state of the art pessimistic approaches. We further develop this research direction and generalize it to the sequential setting, where an online algorithm redeploys the learned policy to the environment and collect more interactions to train the next iteration.
Sequential Off-Policy Learning.
Sequential policy learning extends off-policy learning by allowing the decision-maker to deploy a learned policy, collect additional interactions, and then train the next policy on the entire set of previously gathered data. This approach is widely used in practice since decision systems are continuously updated. Nevertheless, its theoretical foundations remain limited and underexplored, and only a few works have addressed this question.
A related direction is off-policy learning from adaptively collected data, studied in Zhan et al. [2022] and Bibaut et al. [2021]. These works focus on purely offline settings, where data is collected under multiple logging policies with decaying exploration. However, they remain one-shot batch approaches: their goal is to learn a single policy from the collected interactions, without analyzing the effects of redeployment. In contrast, we consider sequential updates and redeployments that occur after the arrival of batches of arbitrary sizes. Our analysis characterizes the conditions under which sequential deployment achieves fast convergence rates. This framework more accurately models real-world scenarios, provides theoretical insight into a widely used practical approach, and unifies the static (fully batch) and dynamic (fully online) regimes. Our proposed algorithms match the convergence rates of the batch setting and further accelerate learning as the behavior policy becomes refined.
Recently, Zenati et al. [2023] extended Counterfactual Risk Minimization [Swaminathan and Joachims, 2015] to the sequential setting, allowing deployments of learned policies within the same framework as ours. Their algorithm achieves accelerated convergence under a Hölderian assumption, but this assumption is not always justified. Moreover, it requires bounded importance weights, which in discrete action spaces is equivalent to assuming uniform coverage. Their method further relies on unrealistic, exponentially growing batch sizes and only uses the most recent batch of data, discarding all past interactions. These limitations arise from the theoretical tools employed, making the algorithm impractical. By contrast, we develop a PAC-Bayesian approach that avoids these shortcomings. Our algorithms are the first to seamlessly interpolate between full batch and full online regimes, achieving the optimal convergence rates of both while requiring only coverage of the optimal actions.
@@
PAC-Bayes learning is a recent branch of learning theory which emerged in the late 90s via the seminal work of [Shawe-Taylor and Williamson, 1997, McAllester, 1998, 1999, 2003] and later pursued by [Catoni, 2003, 2007]. Modern surveys are available to describe the various advances in the field [Guedj, 2019, Alquier, 2024]. PAC-Bayes provide generalization bounds involving a complexity term (often a KL divergence), inspired here from the Bayesian learning paradigm of designing a posterior knowledge of the learning problem based on both training data and a prior knowledge of the considered situation. On the contrary, PAC-Bayes, while inspired from the Bayesian philosophy, relies historically on tools from information theory. This general approach benefits from additional flexibility as PAC-Bayes can be linked and applied to Bayesian learning (see Guedj, 2019) but also blurs the notion of prior and posterior distributions, now independent of the fundamental Bayes formula. This flexibility allowed PAC-Bayes to reach various sub-fields of learning theory: optimization dynamics of learning algorithms [Mou et al., 2018, Andreeva et al., 2024], reinforcement learning [Fard and Pineau, 2010], online learning [Haddouche and Guedj, 2022] and contrastive learning [Nozawa et al., 2020] to name a few.
Appendix B PROOFS OF SECTION 3
In this whole section, we assume that all datasets $S_{k}$ have the same cardinal $n_{k}=m$. This section contains the proof of Sec.˜3.2 and Algorithm˜1.
B.1 Proof of Theorem 3.2
@@ Appendix B PROOFS OF SECTION 3
Plugging this back into Eq.˜1 yields, with probability at least $1-\delta$ over $S$, for any $k\geq 0$, any posterior $Q$:
$\sum_{j=0}^{k}\mathbb{E}_{\theta\sim Q}\left[\sum_{i\in I_{k}}-\log\left(1-\lambda R(d_{\theta})\right)+\log\left(1-\lambda\frac{d_{\theta}(a_{i}\mid x_{i})c_{i}}{\pi_{j}(a_{i}\mid x_{i})}\right)\right]\leq\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right).$
@@
$\log\left(1-\lambda\frac{d_{\theta}(a\mid x)c_{i}}{\pi_{j}(a\mid x)}\right)=d_{\theta}(a,x)\log\left(1-\lambda\frac{c_{i}}{\pi_{j}(a\mid x)}\right).$
@@
B.2 Proof of Proposition 1
We prove the convergence guarantee of Algorithm˜1.
@@ B.2 Proof of Proposition 1
We fix $k\geq 0$, a sequential algorithm Alg and $P$ a data-free prior. To highlight the dependency on $\lambda$, we denote in this proof $\mathcal{S}_{j}=\mathcal{S}_{\lambda,j}$ Our goal is to apply the change of measure inequality on $\mathcal{H}$ to a specific function $f_{k}$. We define this function below, for any sample $S$ and any predictor $\theta$
@@
Plugging this back into Eq.˜2 yields, with probability at least $1-\delta$ over $S$, for any $k\geq 0$, any $Q$:
@@
We also consider $\pi_{k+1}$ the output of Algorithm˜1 time $k$ satisfying:
@@
Then using Sec.˜3.2, and the definition of $\hat{Q}$ yields for any $\lambda$, data-free $P$ with probability at least $1-\delta/2$, for any $k$:
@@
Then, exploiting Eq.˜3 with probability $1-\delta/2$ and taking a union bound yields, with probability at least $1-\delta$, for any $k$: